36. Estimate each of the following, indicating if the actual answer is greater than or less than the estimate.

a. 29/15 + 198/199
b. 5 12/13 + 2 9/10 + 19/20
c. 10 4/9 - 5 1/2 + 99/100

Answer by checkley77(12844)

a. 29/15 + 198/199
Actual: 1.9333+.99498=2.9283
Estimate: 30/15+200/200=2+1=3 estimate is greater.
b. 5 12/13 + 2 9/10 + 19/20
Actual:5.923+2.9=8.823
Estimate:6+3=9 estimate is greater.
c. 10 4/9 - 5 1/2 + 99/100
Actual 10.4-5.5=4.9
Estimate: 10-6=4 estimate is less.